OpenStudy (anonymous):

-(10 - k) - (k - 12) simplify

OpenStudy (anonymous):

-(10 - k) - (k - 12) = -10 + k - k -12 = 1 k - 1 k = 0 -10 - 12 = -10 + -12 = -22 <-- answer